# AgenticTCAD: A LLM-based Multi-Agent Framework for Automated TCAD Code Generation and Device Optimization

### v1

@@ I Introduction
Over the past decades, the semiconductor industry has advanced by following Moore’s law, which has consistently increased transistor density, performance, and energy efficiency[2022_75y]. As CMOS technology approaches its physical and economic limits, however, further scaling is hindered by challenges such as rising leakage currents, variability, and manufacturing costs, making it difficult to sustain historical improvements. To address these issues, the concept of design technology co-optimization (DTCO) has emerged, highlighting the need for joint optimization of process technology and circuit design to continue driving progress in integrated circuits and systems[2024_dtco_ted].
@@
Within the DTCO paradigm, the technology-oriented dimension plays a central role, encompassing material exploration[2023-hi-tcad-material], device architecture innovation[2022-ted-3nm-tcad], and standard-cell layout optimization[2024-iedm-cfet-layout]. Among these tasks, device design and optimization are especially critical, as transistor characteristics directly impact overall system performance. Consequently, Technology Computer-Aided Design (TCAD) simulation[synopsys_tcad] is indispensable for guiding technology scaling and supporting effective DTCO.
@@
TCAD simulation involves highly complex physical mechanisms, such like carrier transport, quantum confinement, and trap effects[2022-iedm-mx2]. Manual device optimization under these conditions is extremely challenging, as the large design space coupled with intricate physics requires significant expertise and effort. In addition, TCAD tools require domain-specific coding languages to describe device structures, meshing, and simulation setups. Since semiconductor device research is often proprietary and lacks open-source references, writing such codes poses a significant barrier for researchers and further limits efficient adoption.
@@
Large language models (LLMs) have recently demonstrated considerable potential in the design domain of DTCO[2025-review-llm4eda]. They have been employed for code generation, enabling the automatic synthesis of design scripts and hardware descriptions directly from natural language inputs[2025-iccad-origen-rtl]. Furthermore, LLMs have been investigated for automated circuit design, offering a pathway from high-level specifications to transistor-level design with reduced human involvement[2024-dac-analogllm].
@@
The technology dimension of DTCO critically depends on device design and optimization. Sentaurus Structure Editor (SDE) and Sentaurus Device (SDevice) are widely used for device construction and electrical simulation, but their domain-specific and proprietary scripts leave the community without publicly accessible datasets for TCAD code generation. Current research efforts have so far focused mainly on SDE script generation. Early attempts explored nanowire device structure generation[2024-nano-tcadllm], but these did not include mesh construction, making direct electrical simulation infeasible. Subsequent work has investigated end-to-end code generation through fine-tuned LLM[2025-chattcad], yet this required translating SDE scripts into Python descriptions beforehand, which introduces additional complexity and has so far been limited to two-dimensional structures. Other approaches have employed agent-based methods to map layout descriptions into SDE scripts[2025-lay2sde], but these often depend on case-specific templates for in-context learning, reflecting the difficulty commercial LLMs still face in handling TCAD-specific syntax.
@@
The remainder of this paper is organized as follows: Section II reviews related work on TCAD simulation, LLM-based code generation, and agent-based methods in EDA. Section III presents the proposed AgenticTCAD framework and its design details. Section IV reports experimental results and Section V concludes the paper. The code and dataset of AgenticTCAD are available at anonymous github.
@@ II-A TCAD Simulation
TCAD simulation has been widely applied in semiconductor device research and development to predict device behavior before fabrication. It supports a variety of material systems and device architectures, including advanced silicon-based technologies [2022-ted-3nm-tcad], III–V semiconductors[2017-ted-tcad35], and emerging devices such as carbon nanotube (CNT)[2019-cnt-tcad]. In addition, TCAD enables comprehensive multiphysics analysis, capturing effects such as defects[2022-iedm-mx2] and self-heating effects[2024-ted-sh-tcad]. Owing to these capabilities, TCAD plays a crucial role in accelerating early-stage device design and optimization.
@@ II-B EDA Code Generation with LLM
Code generation is a key application of large language models (LLMs) and has been widely explored in mainstream programming language[2024-cgen-tranditional]. Within DTCO’s design stage, LLMs have been applied to realize RTL code generation[2025-iccad-origen-rtl, 2025-date-cgen-rl, 2025-date-rtl-cgen], debugging [2024-dac-rlt-fix, 2025-iccad-rtl-de] and optimization [2025-iccad-rewrite]. In the back-end stage, researchers are investigating how LLMs can automatically construct physical design scripts, including flows such as placement and routing, clock-tree synthesis [2024-tcad-chateda]. In the technology domain, there have been initial attempts to use LLMs for generating TCAD SDE input code [2024-nano-tcadllm, 2025-chattcad, 2025-lay2sde] and SDevice decks to support subsequent device simulations, though this line of research remains relatively limited. Our work aims to build a complete TCAD SDE/SDevice code generation flow and release an open-source dataset to foster further investigation in this direction. In summary, LLM-based code generation holds great potential to accelerate the EDA workflow across the entire DTCO iteration process.
@@ II-C Agent Flow in EDA
Agents built upon LLMs provide a flexible paradigm for decomposing complex design tasks into modular, cooperative subtasks. Recent studies have applied LLM-based agents to analog design, such as topology exploration of operational amplifiers[2024-dac-analogllm] and hybrid frameworks that combine LLMs with Bayesian optimization for circuit design[2025-iccad-ADO-LLM]. LLM-based agents have further been investigated in digital design, including recent demonstrations on CPU development with successful tape-out results[2024-dac-chatcpu]. In addition, LLM agents have been explored for generating and optimizing circuit layouts [2025-stdlayout-nv, 2025-tcad-LayoutCopilot], aiming to improve downstream manufacturing quality and design efficiency. Overall, LLM-driven agent flows have shown strong potential to enhance automation and efficiency in EDA.
@@ III-B Domain-Specific LLM Fine-Tuning
After constructing the TCAD-specific dataset, we further adapt a general-purpose LLM to the task of code generation through SFT[2021-sft-openai]. In this stage, the pretrained model is exposed to a large number of carefully designed input–output pairs so that it can gradually adjust its parameters toward the characteristics of TCAD-related tasks. The fine-tuning objective is formulated as a next-token prediction problem, where the model is optimized to maximize the likelihood of generating the correct target sequence conditioned on the given input. Concretely, each training example is represented as a pair: the input corresponds to a natural-language query derived from the prompt templates in Fig. 2(a), while the output is the associated SDE or SDevice script that implements the requested functionality. This design explicitly bridges the gap between high-level textual descriptions and low-level simulation code.
@@ III-C Multi-Agent Workflow for Automated Device Optimization
The workflow in stage 2 executes the TCAD simulation and subsequent post-processing. The SDE script is first executed to construct the NS-FET geometry and corresponding device meshing. The resulting structure is then passed to the SDevice solver, which performs electrical simulations based on the associated physical models, boundary conditions, and solver settings. According to the International Roadmap for Devices and Systems (IRDS)-2024 logic device requirements [IRDS2024], three key performance metrics are selected, and the post-processing module automatically extracts them from the simulated IVI-V characteristics. The on-state current ($I_{on}$) is the drive current that determines both the switching speed of logic gates and the drive strength of the device. The off-state current ($I_{off}$) corresponds to the leakage current in the closed state and directly impacts static power consumption and overall energy efficiency. The subthreshold swing ($SS$) quantifies the sharpness of the switching transition and governs the trade-off between performance and power efficiency. These metrics, together with the complete IVI-V characteristics, are stored in a structured JSON format to enable seamless integration with the subsequent reasoning stage. In addition, the post-processing module also extracs the energy band diagrams in the channel region under different bias conditions. These diagrams include the conduction band ($E_{C}$), the valence band ($E_{V}$), the electron quasi-Fermi level ($E_{Fn}$), and the hole quasi-Fermi level ($E_{Fp}$). Such information provides physical insight into carrier transport and enhances the interpretability of the optimization process.
@@ IV-A Experimental Setup
All experiments were conducted on a Linux-based workstation equipped with two Intel Xeon Platinum 8375C CPUs, each operating at 2.9 GHz with 64 cores, and 4 NVIDIA A6000 GPUs with 48 GB of memory. For SFT, we employed the open-source Qwen2.5-14B-Instruct model[Qwen25], implemented with the Transformers library. The fine-tuning was performed on 30,000 training samples for 5 epochs with a batch size of 8 in bfloat16 precision. The AdamW optimizer was used with momentum parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$, a weight decay of $1\times 10^{-4}$, an initial learning rate of $1\times 10^{-5}$, a cosine learning-rate decay schedule, and a warmup ratio of 0.05. The training procedure required approximately 107 hours on four A6000 GPUs. For TCAD execution, we utilized Synopsys Sentaurus W-2024.09-SP1. The code generation model was deployed on a single A6000 GPU, and the device optimization was carried out using DeepSeek V3.1 reasoning version[deepseekv31].
@@ IV-B Open TCAD Dataset and Model Evaluation
[2024-nano-tcadllm] $\times$ $\times$ $\times$ $\times$
MALTS[2025-lay2sde] $\times$ $\times$ $\times$
ChatTCAD[2025-chattcad] $\times$ $\times$ $\times$
TCADAgent

### v2

@@ I Introduction
Over the past decades, the semiconductor industry has advanced by following Moore’s law, which has consistently increased transistor density, performance, and energy efficiency[13]. As CMOS technology approaches its physical and economic limits, however, further scaling is hindered by challenges such as rising leakage currents, variability, and manufacturing costs, making it difficult to sustain historical improvements. To address these issues, the concept of design technology co-optimization (DTCO) has emerged, highlighting the need for joint optimization of process technology and circuit design to continue driving progress in integrated circuits and systems[11].
@@
Within the DTCO paradigm, the technology-oriented dimension plays a central role, encompassing material exploration[20], device architecture innovation[14], and standard-cell layout optimization[15]. Among these tasks, device design and optimization are especially critical, as transistor characteristics directly impact overall system performance. Consequently, Technology Computer-Aided Design (TCAD) simulation[22] is indispensable for guiding technology scaling and supporting effective DTCO.
@@
TCAD simulation involves highly complex physical mechanisms, such like carrier transport, quantum confinement, and trap effects[19]. Manual device optimization under these conditions is extremely challenging, as the large design space coupled with intricate physics requires significant expertise and effort. In addition, TCAD tools require domain-specific coding languages to describe device structures, meshing, and simulation setups. Since semiconductor device research is often proprietary and lacks open-source references, writing such codes poses a significant barrier for researchers and further limits efficient adoption.
@@
Large language models (LLMs) have recently demonstrated considerable potential in the design domain of DTCO[8]. They have been employed for code generation, enabling the automatic synthesis of design scripts and hardware descriptions directly from natural language inputs[5]. Furthermore, LLMs have been investigated for automated circuit design, offering a pathway from high-level specifications to transistor-level design with reduced human involvement[4].
@@
The technology dimension of DTCO critically depends on device design and optimization. Sentaurus Structure Editor (SDE) and Sentaurus Device (SDevice) are widely used for device construction and electrical simulation, but their domain-specific and proprietary scripts leave the community without publicly accessible datasets for TCAD code generation. Current research efforts have so far focused mainly on SDE script generation. Early attempts explored nanowire device structure generation[17], but these did not include mesh construction, making direct electrical simulation infeasible. Subsequent work has investigated end-to-end code generation through fine-tuned LLM[31], yet this required translating SDE scripts into Python descriptions beforehand, which introduces additional complexity and has so far been limited to two-dimensional structures. Other approaches have employed agent-based methods to map layout descriptions into SDE scripts[7], but these often depend on case-specific templates for in-context learning, reflecting the difficulty commercial LLMs still face in handling TCAD-specific syntax.
@@
The remainder of this paper is organized as follows: Section II reviews related work on TCAD simulation, LLM-based code generation, and agent-based methods in EDA. Section III presents the proposed AgenticTCAD framework and its design details. Section IV reports experimental results and Section V concludes the paper. The code and dataset of AgenticTCAD are available at github.
@@ II-A TCAD Simulation
TCAD simulation has been widely applied in semiconductor device research and development to predict device behavior before fabrication. It supports a variety of material systems and device architectures, including advanced silicon-based technologies [14], III–V semiconductors[1], and emerging devices such as carbon nanotube (CNT)[21]. In addition, TCAD enables comprehensive multiphysics analysis, capturing effects such as defects[19] and self-heating effects[3]. Owing to these capabilities, TCAD plays a crucial role in accelerating early-stage device design and optimization.
@@ II-B EDA Code Generation with LLM
Code generation is a key application of large language models (LLMs) and has been widely explored in mainstream programming language[12]. Within DTCO’s design stage, LLMs have been applied to realize RTL code generation[5, 18, 25], debugging [24, 28] and optimization [29]. In the back-end stage, researchers are investigating how LLMs can automatically construct physical design scripts, including flows such as placement and routing, clock-tree synthesis [27]. In the technology domain, there have been initial attempts to use LLMs for generating TCAD SDE input code [17, 31, 7] and SDevice decks to support subsequent device simulations, though this line of research remains relatively limited. Our work aims to build a complete TCAD SDE/SDevice code generation flow and release an open-source dataset to foster further investigation in this direction. In summary, LLM-based code generation holds great potential to accelerate the EDA workflow across the entire DTCO iteration process.
@@ II-C Agent Flow in EDA
Agents built upon LLMs provide a flexible paradigm for decomposing complex design tasks into modular, cooperative subtasks. Recent studies have applied LLM-based agents to analog design, such as topology exploration of operational amplifiers[4] and hybrid frameworks that combine LLMs with Bayesian optimization for circuit design[30]. LLM-based agents have further been investigated in digital design, including recent demonstrations on CPU development with successful tape-out results[26]. In addition, LLM agents have been explored for generating and optimizing circuit layouts [9, 16], aiming to improve downstream manufacturing quality and design efficiency. Overall, LLM-driven agent flows have shown strong potential to enhance automation and efficiency in EDA.
@@ III-B Domain-Specific LLM Fine-Tuning
After constructing the TCAD-specific dataset, we further adapt a general-purpose LLM to the task of code generation through SFT[2]. In this stage, the pretrained model is exposed to a large number of carefully designed input–output pairs so that it can gradually adjust its parameters toward the characteristics of TCAD-related tasks. The fine-tuning objective is formulated as a next-token prediction problem, where the model is optimized to maximize the likelihood of generating the correct target sequence conditioned on the given input. Concretely, each training example is represented as a pair: the input corresponds to a natural-language query derived from the prompt templates in Fig. 2(a), while the output is the associated SDE or SDevice script that implements the requested functionality. This design explicitly bridges the gap between high-level textual descriptions and low-level simulation code.
@@ III-C Multi-Agent Workflow for Automated Device Optimization
The workflow in stage 2 executes the TCAD simulation and subsequent post-processing. The SDE script is first executed to construct the NS-FET geometry and corresponding device meshing. The resulting structure is then passed to the SDevice solver, which performs electrical simulations based on the associated physical models, boundary conditions, and solver settings. According to the International Roadmap for Devices and Systems (IRDS)-2024 logic device requirements [10], three key performance metrics are selected, and the post-processing module automatically extracts them from the simulated IVI-V characteristics. The on-state current ($I_{on}$) is the drive current that determines both the switching speed of logic gates and the drive strength of the device. The off-state current ($I_{off}$) corresponds to the leakage current in the closed state and directly impacts static power consumption and overall energy efficiency. The subthreshold swing ($SS$) quantifies the sharpness of the switching transition and governs the trade-off between performance and power efficiency. These metrics, together with the complete IVI-V characteristics, are stored in a structured JSON format to enable seamless integration with the subsequent reasoning stage. In addition, the post-processing module also extracs the energy band diagrams in the channel region under different bias conditions. These diagrams include the conduction band ($E_{C}$), the valence band ($E_{V}$), the electron quasi-Fermi level ($E_{Fn}$), and the hole quasi-Fermi level ($E_{Fp}$). Such information provides physical insight into carrier transport and enhances the interpretability of the optimization process.
@@ IV-A Experimental Setup
All experiments were conducted on a Linux-based workstation equipped with two Intel Xeon Platinum 8375C CPUs, each operating at 2.9 GHz with 64 cores, and 4 NVIDIA A6000 GPUs with 48 GB of memory. For SFT, we employed the open-source Qwen2.5-14B-Instruct model[23], implemented with the Transformers library. The fine-tuning was performed on 30,000 training samples for 5 epochs with a batch size of 8 in bfloat16 precision. The AdamW optimizer was used with momentum parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$, a weight decay of $1\times 10^{-4}$, an initial learning rate of $1\times 10^{-5}$, a cosine learning-rate decay schedule, and a warmup ratio of 0.05. The training procedure required approximately 107 hours on four A6000 GPUs. For TCAD execution, we utilized Synopsys Sentaurus W-2024.09-SP1. The code generation model was deployed on a single A6000 GPU, and the device optimization was carried out using DeepSeek V3.1 reasoning version[6].
@@ IV-B Open TCAD Dataset and Model Evaluation
[17] $\times$ $\times$ $\times$ $\times$
MALTS[7] $\times$ $\times$ $\times$
ChatTCAD[31] $\times$ $\times$ $\times$
AgenticTCAD